OpenStudy (anonymous):

Solve. x2 – x = 0 x = 0 x = 1 and x = –1 x = 0 and x = –1 x = 0 and x = 1

OpenStudy (anonymous):

x^2 -x = 0 x (x-1) = 0 x = 0 x = 1

jimthompson5910 (jim_thompson5910):

x^2-x = 0 x(x-1) = 0 x = 0 or x-1=0 x=0 or x=1

jimthompson5910 (jim_thompson5910):

So the answer is choice D

OpenStudy (anonymous):

fourth one